OpenStudy (anonymous):

Verify the identity (1+ sinx)/(cos x) + (cos x)/(1+sin x) = 2 sec x?

OpenStudy (anonymous):

@goformit100 can you help?

OpenStudy (goformit100):

Are you Indian ?

OpenStudy (anonymous):

no why?

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

o.O

OpenStudy (goformit100):

I know the answer

OpenStudy (anonymous):

Right, hold on @pvs285

OpenStudy (goformit100):

(1+ sinx)/(cos x) = (1/cos x + sinx/cos x) = ?

OpenStudy (anonymous):

(1/cosx + tanx) right

OpenStudy (goformit100):

yes Now do same for (cos x)/(1+sin x)

OpenStudy (goformit100):

What did you got ?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

is it cosx+tanx/1?

OpenStudy (anonymous):

@goformit100 are you still here?

OpenStudy (goformit100):

yes now add them up

OpenStudy (anonymous):

(1/cosx+tanx)+(cosx+tanx/1)

myininaya (myininaya):

My first step would have been to combine the fractions.

OpenStudy (anonymous):

(cosx+tanx)(cosx+tanx)/cosx+tanx

OpenStudy (anonymous):

what now

OpenStudy (anonymous):

plz help

myininaya (myininaya):

Yeah I don't like all those tangents. I still would have approached it by combining the fractions as a first step.

OpenStudy (anonymous):

how do i do that

myininaya (myininaya):

Force a common denominator by doing: a/b+c/d =(ad+bc)/(bd)

myininaya (myininaya):

\[\frac{a}{b}+\frac{c}{d}=\frac{ad+cb}{bd}\]

OpenStudy (anonymous):

whats a b c and d

myininaya (myininaya):

a is the numerator of the first fraction b is the denominator of the first fraction c is the numerator of the second fraction d is the denominator of the second fraction

OpenStudy (anonymous):

(cosx*cosx)+(1+sin*1+sin)/(1+sin*cosx)

OpenStudy (anonymous):

Or you could write it as \[=\frac{1+s}{c}+\frac{c}{1+s} \\ =(\frac{1+s}{c}\times \frac{1+s}{1+s})+(\frac{c}{1+s} \times \frac{c}{c})\] \[=\frac{1+2s+s^2+c^2}{c+cs}\]

myininaya (myininaya):

I think you wrote: \[\frac{\cos(x) \cos(x)+(1+\sin(x))(1+\sin(x))}{(1+\sin(x))(\cos(x))}\] correct?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

what now

myininaya (myininaya):

Now multiply (1+sin(x))(1+sin(x)) out

OpenStudy (anonymous):

1+2sinx

myininaya (myininaya):

+what else

OpenStudy (anonymous):

what do i do now

myininaya (myininaya):

You haven't multiplied (1+sin(x))(1+sin(x)) correctly yet

OpenStudy (anonymous):

cos^2x+1+2sinx

OpenStudy (anonymous):

is that correct for the numerator

myininaya (myininaya):